Calculate and categorize the Point elasticity for the (inverse)demand curve P=24-2Q at p1=14,p2=12,p3=10
Answer #1

The inverse demand curve can be rewritten as

2Q = 24-P

or Q = 12-0.5P

so, dQ/dP = -0.5

Elasticity formula = dQ/dP*P/Q

A) When P=14

Q = 12-0.5*14 = 5

E = -0.5*14/5 = -1.4

B) When P=12

Q = 12-0.5*12 = 6

E = -0.5*12/6 = -1

C) When P = 10

Q = 12-0.5*10 = 7

E = -0.5*10/7 = -0.71